OpenStudy (anonymous):

a box of chocolates contains 12 chocolates with nuts and 24 24 with out nuts. Determine the probability of choosing 2 chocolates with nuts in a row without replacement

OpenStudy (anonymous):

\[\frac{12}{36}*\frac{11}{35}\\ \frac{1}{3}*\frac{11}{35}\\ =\frac{11}{105}\]